OpenStudy (anonymous):

differentiatef(x)=1/1+x^2

OpenStudy (bahrom7893):

I suck at quotient rule so I'll use the power and chain rules.. f(x) = (1+x^2)^(-1) f'(x) = -1(1+x^2)^(-2)*(2x)

OpenStudy (bahrom7893):

or f'(x) -2x/(1+x^2)^2

OpenStudy (anonymous):

2x

OpenStudy (bahrom7893):

ackhat, i dont think he meant that

OpenStudy (anonymous):

))))

OpenStudy (bahrom7893):

but i agree, people do need to learn how to use parenthesis

OpenStudy (anonymous):

(1 + x^2)' = 2x f'(x) -2x/(1+x^2)^2